40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 13 00:56:07 2016, gautam, Update, endtable upgrade, X endtable repopulation 
    Reply  Wed Apr 13 11:00:28 2016, Steve, Update, endtable upgrade, X endtable repopulation 
       Reply  Wed Apr 13 18:25:07 2016, gautam, Update, endtable upgrade, Lightwave health check 
          Reply  Fri Apr 15 18:38:12 2016, gautam, Update, endtable upgrade, Lightwave health check - NO IMPROVEMENT 
             Reply  Tue Apr 19 18:37:29 2016, gautam, Update, endtable upgrade, Laser swap + optical layout IR_modematch_19April2016_2.pdf
                Reply  Thu Apr 21 14:25:52 2016, gautam, Update, endtable upgrade, Green light recovered IMG_6567.JPGIMG_6568.JPGCVI_reflectivity.jpeg
                Reply  Tue Apr 26 15:22:35 2016, Steve, Update, endtable upgrade, Cleaning ETMX vacuum dirty window ETMX-Tvp.jpgETMX-TvpDetail.jpg
                   Reply  Tue Apr 26 23:19:42 2016, gautam, Update, endtable upgrade, Green aligned to arm - high order mode flashes seen IMG_2219.JPGIMG_2220.JPG
                      Reply  Thu Apr 28 00:41:08 2016, gautam, Update, endtable upgrade, more progress - Transmon PD installed 
                         Reply  Fri Apr 29 00:55:46 2016, gautam, Update, endtable upgrade, green PDH locked to Xarm 
                            Reply  Mon May 2 19:14:18 2016, gautam, Update, endtable upgrade, Optical layout almost complete ETMX_wedge.pdf
                               Reply  Thu May 5 03:05:37 2016, gautam, Update, endtable upgrade, ALS status update ALS_noiseSpec_5May2016_2.pdfCoherence_5May2016.pdfimage.jpeg
                                  Reply  Thu May 5 04:05:03 2016, ericq, Update, LSC, Aux X PDH checks nom_noise.pngnom_olg.png
                                     Reply  Thu May 5 14:03:52 2016, ericq, Update, LSC, Further Aux X PDH tweaks 2016-05-05_newals.pdf
                                  Reply  Thu May 5 14:05:01 2016, rana, Update, endtable upgrade, ALS status update 
                                  Reply  Thu May 5 21:28:44 2016, gautam, Update, endtable upgrade, Innolight PZT capacitance 
                               Reply  Thu May 19 16:29:20 2016, Steve, Update, endtable upgrade, Optical layout almost complete ETMX_4x3_closed.jpgsealedETMXenclosure.jpg
                                  Reply  Fri Sep 30 19:53:07 2016, gautam, Update, endtable upgrade, X end IR pickoff fiber coupled 
                                     Reply  Wed Oct 5 16:28:10 2016, gautam, Update, endtable upgrade, EX laser power monitor PD installed 
                   Reply  Wed Apr 27 15:04:47 2016, Steve, Update, endtable upgrade, Cleaning ETMX vacuum dirty window 1cETMX-Tcmp.jpg
                      Reply  Fri Apr 29 16:05:23 2016, gautam, Update, endtable upgrade, Cleaning ETMX vacuum dirty window IMG_0755.JPG
Message ID: 12090     Entry time: Tue Apr 26 23:19:42 2016     In reply to: 12089     Reply to this: 12095
Author: gautam 
Type: Update 
Category: endtable upgrade 
Subject: Green aligned to arm - high order mode flashes seen 

Attachment #1

Layout as of today. Most of the green path is done. The Green REFL PD + PZT mirrors have not been hooked up to their respective power sources yet (I wonder if it's okay to start laying cables through the feedthroughs on either end of the table already, or if we want to put whatever it is that makes it airtight eventually in first?). A rough power budget has been included (with no harmonic separator just before the window), though some optimization can be done once the table is completely repopulated.

Attachment #2

A zoomed-in version of the REFL path.

Some general notes:

  1. I've tried to use the custom 3/4" O.D. posts + baseplate arrangement wherever possible (only 1 steering mirror is on a 1" post clamped with a fork to the table because of space constraints). Where the baseplates could not be bolted onto the table directly, I've used Newport SS Dogs to do the job.
  2. I checked for continuity between the PZT outer case and the table top with a multimeter, and found none. So I chose to leave the Thorlabs baseplates in place. For the REFL PD, I've used an insulating baseplate given to me by Steve.
  3. I've used some custom length 3/4" O.D. posts to get the beam up to the right height (~4.75") just before sending the green beam in. The beam height is 4" elsewhere.
  4. I was playing around with positioning the harmonic separator immediately before the vacuum chamber window - I found that there is a substantial amount of green light that is reflected, though there doesn't seem to be any IR leaking through. The mirror was labelled Y1-1037-45P, which is a code for CVI mirrors, though I believe it is a LaserOptik product and that we have a couple of other such mirrors in the optics cabinet - though they are all 1". This document suggests that from the back side, there should be <0.1% reflection of green while on the front side it should be < 3%. I will have to hunt a little more for the specs, and measure the powers to see if they match the previously quoted numbers. In any case, I'll have to think of how to separate the (unwanted) reflected green and the transmitted IR from the cavity in the IR transmon path.
  5. There are some minor changes to the planned layout posted here - I will update these in due course once the Transmon path and Oplev have been set up.

I am closing the PSL shutter and the EX laser shutters for the night as I have applied a layer of first contact to the window for cleaning purposes, and we don't want any laser light incident on it. It may be that the window is so dirty that we may need multiple F.C. cleaning rounds, we will see how the window looks tomorrow...

 

Attachment 1: IMG_2219.JPG  5.250 MB  | Hide | Hide all | Show all
IMG_2219.JPG
Attachment 2: IMG_2220.JPG  5.872 MB  | Show | Hide all | Show all
ELOG V3.1.3-